OpenStudy (anonymous):

A train goes up a hill with a 15º incline. If the train has constant speed of 22 m/s, what are the vertical and horizontal components of the train's velocity? (Remember to use the law of sine and cosines for this problem)

OpenStudy (anonymous):

 rmb the sine and cosine: \[\sin \theta=\frac{ v_{y} }{ \left| v \right| }\] \[\cos\theta=\frac{ v_{x} }{ \left| v \right| }\]